### 17 Jan 2024 v1

# High Confidence Level Inference is Almost Free using Parallel Stochastic Optimization

Abstract
Uncertainty quantification for estimation through stochastic optimization solutions in an online setting has gained popularity recently. This paper introduces a novel inference method focused on constructing confidence intervals with efficient computation and fast convergence to the nominal level. Specifically, we propose to use a small number of independent multi-runs to acquire distribution information and construct a $t$-based confidence interval. Our method requires minimal additional computation and memory beyond the standard updating of estimates, making the inference process almost cost-free. We provide a rigorous theoretical guarantee for the confidence interval, demonstrating that the coverage is approximately exact with an explicit convergence rate and allowing for high confidence level inference. In particular, a new Gaussian approximation result is developed for the online estimators to characterize the coverage properties of our confidence intervals in terms of relative errors. Additionally, our method also allows for leveraging parallel computing to further accelerate calculations using multiple cores. It is easy to implement and can be integrated with existing stochastic algorithms without the need for complicated modifications.
1 Introduction
Consider the statistical inference problem for model parameters where the true model parameter $x^{*}\in\mathbb{R}^{d}$ can be characterized as the minimizer of an objective function $F(x)$ from $\mathbb{R}^{d}$ to $\mathbb{R}$, i.e,
$x^{*}=\underset{x\in\mathbb{R}^{d}}{\arg\min}\ F(x).$ (1)
The objective function $F(x)$ is defined as $F(x)=\mathbb{E}_{\xi\sim\Pi}{f(x,\xi)}$, where $f(x,\xi)$ represents a noisy measurement of $F(x)$, and $\xi$ is a random variable following the distribution $\Pi$.
As modern datasets grow increasingly large and are often collected in an online manner, classical deterministic optimization which requires storing all data is becoming less appealing or even infeasible for such problems. One solution is to employ stochastic approximation (SA), updating the estimate of the minimizer based on the stochastic gradient/subgradient. This approach is characterized by low memory and computational efficiency, making it suitable for an online setting with sequential data and decision-making. One of the most popular algorithms is the stochastic gradient descent (SGD), also known as the Robbins-Monro algorithm [30]. A diverse range of variants have been developed to accelerate convergence or reduce variance in different scenarios [31, 27, 7, 28, 14, 34, 37]. Beyond convergence analysis of the final (point) estimate, there is a growing emphasis on uncertainty quantification and statistical inference within the context of SA algorithms, which forms a significant part of recent research endeavors [4, 10, 17, 33].
With streaming data $\{\xi_{i}\}_{i\geq 1}$, and assuming we obtain iterates/outputs of an SA algorithm, the primary goal of this paper is to enhance statistical inference by constructing confidence intervals based on these iterates in an online setting. Specifically, for a given vector $\upsilon\in\mathbb{R}^{d}$, we aim to construct a valid $(1-\alpha)\times 100\%$ confidence interval $\widehat{\textnormal{CI}}$ for the linear functional $\upsilon^{\top}x^{*}$, that is
$\mathbb{P}(\upsilon^{\top}x^{*}\in\widehat{\textnormal{CI}})-(1-\alpha)=\alpha-\mathbb{P}(\upsilon^{\top}x^{*}\not\in\widehat{\textnormal{CI}})\approx 0,$ (2)
where $\alpha\in(0,1)$. To fit in an online setting, the proposed confidence interval can be updated recursively as new data becomes available. It utilize only previous SA iterates, requiring minimal extra computation for the inference purpose beyond the original computation, thus allowing for easy integration into existing codebases.
In this paper, we consider a high level of confidence, i.e., $\alpha\approx 0$, as uncertainty quantification is particularly important in applications involving high-stakes decisions, where a nearly 100% confidence interval is required. Moreover, with datasets growing increasingly large, the demand for higher-level confidence intervals becomes more prevalent. Additionally, in applications involving multiple simultaneous tests, such as high-dimensional parameter analysis, correction techniques like the Bonferroni method are employed. This leads to each individual test maintaining a sufficiently high confidence level (related to dimension). In such cases, the guarantee in (2) may not be sufficient. In particular, we shall construct confidence intervals $\widehat{\textnormal{CI}}$ such that the relative error
$\Delta_{\alpha}:=\left|\frac{\mathbb{P}(\upsilon^{\top}x^{*}\in\widehat{\textnormal{CI}})-(1-\alpha)}{\alpha}\right|=\left|\frac{\mathbb{P}(\upsilon^{\top}x^{*}\not\in\widehat{\textnormal{CI}})}{\alpha}-1\right|$ (3)
is small. Note that (3) offers a much more refined assessment than (2). For example, if $\alpha=10^{-4}$ and $\mathbb{P}(\upsilon^{\top}x^{*}\not\in\widehat{\textnormal{CI}})=10^{-3}$, then (2) is not severely violated, while $\Delta_{\alpha}$ in (3) is very different from 00. In this context, it is crucial to recognize that even a slight undercoverage can be significant due to low tolerance for error. Conversely, an extremely wide confidence interval that nearly always covers can become uninformative, underscoring the importance of precision in interval construction. Hence, employing a method that provides confidence intervals with rapid convergence to the desired coverage level is essential. In Section 3, we derive the upper bound (with explicit rate) of the relative error of coverage for the constructed confidence intervals and explicitly detail the dependence on $\alpha$ in the upper bound. The results indicate that our method remains valid even when $\alpha$ is potentially very small or decreases with the total sample size or the number of hypotheses.
1.1 Background: existing confidence interval construction
Practical inference methods are based on the limiting distribution of SA solutions. Consider the vanilla SGD iterates with the recursion form:
$x_{i}=x_{i-1}-\eta_{i}\nabla f(x_{i-1},\xi_{i}),\quad i=1,2,\ldots,$
where $\nabla{f}(x,\xi)$ is the gradient vector of $f(x,\xi)$ with respect to the first variable $x$, and $\eta_{i}$ is the step size at the $i$-th step. Denoted by $\bar{x}_{n}=n^{-1}\sum_{i=1}^{n}x_{i}$ the average of all past iterations. In the celebrated work of [27], it is shown that, under suitable conditions, the averaged SGD (ASGD) exhibits the asymptotic normality, that is,
$\sqrt{n}(\bar{x}_{n}-x^{*})\Rightarrow\mathcal{N}(0,\Sigma),$ (4)
where $\Sigma=A^{-1}SA^{-1}$ is the sandwich form covariance matrix with $A=\nabla^{2}F(x^{*})$ and $S=\mathbb{E}\left([\nabla f(x^{*},\xi)][\nabla f(x^{*},\xi)]^{T}\right).$ Note that the asymptotic normality result is the same as that for offline M-estimators [35], and achieves asymptotic minimax optimality [12, 8]. Similar asymptotic normality results have been established for other variants of SGD with adjusted asymptotic covariance matrices [18, 36, 26].
These asymptotic normality results form the foundation of statistical inference in an online setting. As the limiting covariance matrix is unknown in practice, to perform practical inference, there are three primary methods for constructing confidence intervals.
  • The first method relies on recursively estimating the asymptotic covariance matrix $\Sigma$. [4] proposes the plug-in method to estimate $A$ and $S$ separately using sample averages and then applying them in the sandwich form. [40] proposed the online batch-means method, which only utilizes SGD iterates and is more computationally efficient. Both methods provide consistent estimators for the asymptotic covariance matrix $\Sigma$ of ASGD solutions. With a consistent covariance estimate $\widehat{\Sigma}_{n}$, one can construct confidence intervals for $\upsilon^{\top}x^{*}$ as
    $\widehat{\textnormal{CI}}_{n,\textnormal{cov}}=\left[\upsilon^{\top}\bar{x}_{n}-z_{1-\alpha/2}\sqrt{\frac{\upsilon^{\top}\widehat{\Sigma}_{n}\upsilon}{n}},\ \upsilon^{\top}\bar{x}_{n}+z_{1-\alpha/2}\sqrt{\frac{\upsilon^{\top}\widehat{\Sigma}_{n}\upsilon}{n}}\right],$
    where $z_{1-\alpha/2}$ is the $(1-\alpha/2)\times 100\%$ percentile of the standard normal distribution.
  • The second method takes advantage of statistical pivotal statistics. One example is the random scaling method. Instead of consistently estimating the asymptotic covariance matrix, [17] leverages the asymptotic normality result by constructing asymptotic pivotal statistics after self-normalization. Specifically, they studentize $\sqrt{n}(\bar{x}_{n}-x^{*})$ via the random scaling matrix
    $\widehat{V}_{rs,n}=\frac{1}{n}\sum_{s=1}^{n}\left\{\frac{1}{\sqrt{n}}\sum_{i=1}^{s}(x_{i}-\bar{x}_{n})\right\}\left\{\frac{1}{\sqrt{n}}\sum_{i=1}^{s}(x_{i}-\bar{x}_{n})\right\}^{T}.$
    The resulting statistic is asymptotically pivotal and the confidence interval for $\upsilon^{\top}x^{*}$ is then constructed as
    $\widehat{\textnormal{CI}}_{n,\textnormal{rs}}=\left[\upsilon^{\top}\bar{x}_{n}-q_{rs,1-\alpha/2}\sqrt{\frac{\upsilon^{\top}\widehat{V}_{rs,n}\upsilon}{n}},\ \upsilon^{\top}\bar{x}_{n}+q_{rs,1-\alpha/2}\sqrt{\frac{\upsilon^{\top}\widehat{V}_{rs,n}\upsilon}{n}}\right],$ (5)
[…]
  • An alternative method for inference is via bootstrap. One can apply bootstrap perturbations and modify the original SGD path. Then, the asymptotic distribution of the online estimate as well as other quantities such as variance or quantiles can be estimated using a large number of bootstrapped sequences [10, 19, 33].
Same ideas have also been applied for inference when using different algorithms or dealing with online decision-making problems [23, 21, 3, 29, 33]. Note that all the three methods above have their advantages and applicable use cases. The first and third methods can provide consistent estimators of the limiting covariance matrix. However, the cost of using bootstrap (the third method) involves heavy computation or complicated modification to existing code base, and we will not consider this method. The online covariance matrix estimation (the second method) is a difficult task in SGD settings. The plug-in estimator requires Hessian information which is typically unavailable, and involves matrix computation that requires an $O(d^{3})$ computational cost, which is not desirable for large dimensions. The online batch-means methods do not require extra information such as the Hessian and are computationally efficient, but they come at the cost of slow convergence. The random scaling method does not provide a consistent covariance matrix estimator, yet in terms of confidence interval construction, it is computationally comparable to the online batch-means method while offering better coverage. However, the critical values of the self-normalized statistics are not easy to obtain for arbitrary $\alpha$, we simulate the value via MCMC in Section 4.
In terms of theoretical guarantees, although all three methods demonstrate asymptotically valid coverage of confidence intervals, the convergence guarantee without a specific rate and without explicit dependence on $\alpha$ is not sufficient, as demonstrated above and in Section 3. This limitation is not significant at moderate confidence levels but becomes substantial at higher levels, leading to unstable coverage. This may result in either undercoverage (failing to meet the standard) or overcoverage (producing an excessively wide confidence interval, thereby diminishing the interval’s meaningfulness). A detailed comparison and discussion of these methods can be found in [17]. We also make a brief summary in Table 1 comparing the above online inference methods.
Method Plug-In Online BM Random Scale This paper
consistent covariance estimator? / /
to avoid Hessian? /
CI coverage convergence rate? / / /
[…]
Table 1: Comparison of methods for online statistical inference method: Plug-In [4], Online BM [40], Random Scale [17].
1.2 Contribution
We propose a new method that utilizes a small number of parallel runs to effectively acquire information about the distribution, while maintaining a fully online status. We demonstrate that the confidence intervals constructed via the parallel inference method provide asymptotically exact coverage with more rigorous theoretical guarantees than existing methods, featuring an explicit convergence rate of the relative error of coverage. Additionally, this method offers better coverage compared to other methods, as demonstrated in Section 4. It is also the most computationally efficient among all considered inference methods. Our approach avoids the heavy cost of resampling. Unlike methods based on covariance matrix estimation or the random scaling method, our method does not require updating a $d\times d$ matrix at each iteration. Additional computation or memory for inference beyond running SGD is required only when necessary at specific steps and is minimal, making the inference almost free. Another advantage of our method is its suitability for settings where parallel computing is needed, which can further accelerate computation. This is particularly relevant in scenarios such as processing extremely large and high-frequency datasets, or in federated learning scenarios where data are distributed across different clients [41, 6, 20, 13, 24, 11]. In our work, the requirement for parallel processing is seen not as a burden but as a beneficial tool.
2 Inference with parallel runs of stochastic algorithms
In this section, we introduce the parallel run inference method for constructing confidence intervals. The method involves $K$ parallel runs of a predetermined stochastic algorithm, calculating the sample variance of the linear functional of interest from $K$ parallel runs, and self-normalizing to obtain asymptotic pivotal $t$-statistics and the corresponding confidence interval.
2.1 Parallel computing
Consider a general stochastic algorithm characterized by the update rule $h_{i}$ at the $i$-th step and $K$ parallel run sequences. For each of the $k$-th sequence where $k=1,\ldots,K$, we begin with a random initialization $\widehat{x}_{0}^{(k)}$. The estimate for the $k$-th machine at the $i$-th iterate is denoted by $\widehat{x}_{i}^{(k)}$. The recursive update is given by
$\widehat{x}_{i}^{(k)}=h_{i}(\xi_{i}^{(k)},\mathcal{F}_{i-1}^{(k)}),\quad i=1,2,\ldots,$ (6)
where $\mathcal{F}_{i-1}^{(k)}=\sigma(\xi_{i-1}^{(k)},\xi_{i-2}^{(k)},\ldots)$ encapsulates information from the previous step, such as $\widehat{x}_{i-1}^{(k)}$ or other intermediate estimates according to the algorithm. For example, in the case of ASGD, we have
$\begin{cases}x_{i}^{(k)}=x_{i-1}^{(k)}-\eta_{i}\nabla f(x_{i-1}^{(k)},\xi_{i}^{(k)}),\\ \widehat{x}_{i}^{(k)}=\{(i-1)\widehat{x}_{i-1}^{(k)}+\widehat{x}_{i}^{(k)}\}/i,\end{cases}$ (7)
where $\nabla f(x_{i-1}^{(k)},\xi_{i}^{(k)})$ is the derivative of the objective function $f$ with respect to the first variable, and step size is usually chosen as $\eta_{i}=\eta\times i^{-\beta}$ for some $\beta\in(0.5,1]$. If we seek output for estimation or inference at the $n$-th step (with $N=nK$ total samples), we can aggregate the results from the $K$ machines by averaging the estimates or predictions. Specifically, we define the parallel average of the $K$ sequences as
$\bar{x}_{K,n}=\frac{1}{K}\sum_{k=1}^{K}\widehat{x}_{n}^{(k)}.$ (8)
In a practical online setting, sequential data $\{\xi_{i}\}_{i=1,2,...}$ can be distributed across $K$ different machines, with $\xi_{i}^{(k)}=\xi_{k+K(i-1)}$. Alternatively, in an offline setting, the dataset can be randomly divided into $K$ batches. Note that when the initialization $\widehat{x}_{0}^{(k)}$ is the same for all $k=1,\ldots,K$, the output from $K$ sequences $\widehat{x}_{i}^{(k)}$, $k=1,\ldots,K$, will be independent and identically distributed (i.i.d.), given that the data components $\xi_{i}^{(k)}$ are i.i.d..
Note that, unlike local/federated SGD as discussed in [39, 21, 37], we do not communicate local solutions/gradients to obtain a common averaged iterate for parallel runs at intermediate iterations. Our method is more akin to model averaging, often referred to as one-shot averaging [41]. On one hand, model averaging without communication costs can still achieve good convergence when $K$ is small or moderate. On the other hand, it ensures the $K$ sequences are i.i.d. and enables us to construct a asymptotically pivotal $t$-statistic and demonstrate strong convergence in a later section. Additionally, the straightforward parallel running and model averaging make it easier to apply to different stochastic algorithms. In some cases, local updates with more frequent periodic averaging would improve statistical efficiency of the algorithm and communication costs may not be a problem. The inference procedure may still hold with refined proof. However, discussing the difference between vanilla SGD, parallel SGD and local SGD is beyond the scope of this paper.
2.2 Asymptotic $t$-distribution
In this context, various stochastic approximation algorithms can be employed to run parallel sequences. To derive a valid $t$-distribution, it is essential to consider cases where asymptotic normality is applicable for the estimate $\widehat{x}_{n}^{(k)}$ in each sequence. Specifically, for each $k=1,\ldots,K$,
$\sqrt{n}(\widehat{x}_{n}^{(k)}-x^{*})\Rightarrow\mathcal{N}(0,\Sigma),\enspace\mathrm{as}\enspace n\to\infty.$
In the case of ASGD as denoted in (7), the celebrate work of Polyak and Juditsky [27] demonstrated the asymptotic normality with the sandwich form $\Sigma$ as mentioned before. Other algorithms, such as various versions of weighted-averaged SGD [36], Root-SGD [18], and StoSQP [26], have also been shown to possess this asymptotic normality property, albeit with adjusted limiting covariance matrices.
For any vector $\upsilon\in\mathbb{R}^{d}$, considering inference for the linear functional $\upsilon^{\top}x^{*}$ at the $n$-th iteration (with $N=nK$ total samples), define the sample variance $\widehat{\sigma}_{\upsilon}^{2}$ as
$\widehat{\sigma}_{\upsilon}^{2}=\frac{1}{K-1}\sum_{k=1}^{K}(\upsilon^{\top}\widehat{x}_{n}^{(k)}-\upsilon^{\top}\bar{x}_{K,n})^{2},$
where $\bar{x}_{K,n}$ is the sample average defined in (8). It is worth noting that $\widehat{\sigma}_{\upsilon}^{2}$ is not a consistent estimator for the variance of $\upsilon^{\top}x^{*}$. However, we can studentize $\sqrt{K}(\upsilon^{\top}\bar{x}_{K,n}-\upsilon^{\top}x^{*})$ with $\widehat{\sigma}_{\upsilon}$ to obtain a $t$-statistic which is asymptotically pivotal. Assuming the validity of the asymptotic normality result, together with the i.i.d. property of $\{\widehat{x}_{n}^{(k)}\}_{k=1,...,K}$, we can infer a $t$-type distribution, that is,
$\widehat{t}_{\upsilon}:=\frac{\sqrt{K}(\upsilon^{\top}\bar{x}_{K,n}-\upsilon^{\top}x^{*})}{\widehat{\sigma}_{\upsilon}}\Rightarrow t_{K-1}.$ (9)
Based on (9), we can construct a $(1-\alpha)\times 100\%$ confidence interval for $\upsilon^{\top}x^{*}$ as follows,
$\widehat{\textnormal{CI}}_{\upsilon}=\left[\upsilon^{\top}\bar{x}_{K,n}-\frac{t_{1-\alpha/2,K-1}\widehat{\sigma}_{\upsilon}}{\sqrt{K}},\enspace\upsilon^{\top}\bar{x}_{K,n}+\frac{t_{1-\alpha/2,K-1}\widehat{\sigma}_{\upsilon}}{\sqrt{K}}\right],$ (10)
where $t_{1-\alpha/2,K-1}$ is the $(1-\alpha/2)\times 100\%$ percentile for the $t_{K-1}$ distribution. The proposed confidence interval in (10) is fairly easy and efficient to construct. In particular, when $K=2$, $\widehat{t}_{\upsilon}$ is asymptotically distributed as the standard Cauchy distribution with upper quantile $t_{1-\alpha/2,1}=\tan((1-\alpha)\pi/2)$. The entire procedure is summarized in Algorithm 1. For an illustration, refer to Figure 1.
Remark 1 (Almost cost-free).
We observe that the inference step in our method can be performed whenever necessary with minimal calculation and memory requirements, without needing any modifications to existing stochastic algorithms. This makes it almost cost-free and can be easily integrated into existing codebases. In contrast, all other methods typically demand considerable extra effort for inference. This may involve complex modifications, as seen in [33], or entail storing and updating a matrix at each iteration, as required by covariance-matrix-estimation-based methods or the random scaling method. In these cases, the computing and memory costs for inference purposes usually far exceed those involved in the SGD update itself.
Remark 2 (Choice of $K$).
A larger $K$ brings more stable and usually shorter (in expectation) confidence intervals, as indicated in (10). And a smaller $K$ will decrease the effect of sample splitting and lead to better convergence of the estimate in a single run. As demonstrated in Section 4, the performance of inference is not that sensitive to the choice of $K$ when $K$ is in a reasonable range, and $K=6$ is a good choice in practice. Users can opt for a smaller $K$ when dealing with a moderate dataset to ensure sufficient sample size and faster convergence in a single trial. If parallel computing resources are available and preferred, especially when dealing with very large datasets or high data acquisition rates, users are encouraged to utilize more machines.
Input: stochastic algorithm $h$, number of parallel runs $K$
for $i=1,2,\ldots$ do
       for $k=1,\ldots,K$ do
             Update $\widehat{x}_{i}^{(k)}=h_{i}(\xi_{i}^{(k)},\mathcal{F}_{i-1}^{(k)})$ ($\xi_{i}^{(k)}$ is data received);
       end for
      Output if necessary:
       $\bar{x}_{K,i}\leftarrow\frac{1}{K}\sum_{k=1}^{K}\widehat{x}_{k,i}$;
       $\widehat{\sigma}_{\upsilon}^{2}\leftarrow\frac{1}{K-1}\sum_{k=1}^{K}\left(\upsilon^{\top}\widehat{x}_{k,i}-\upsilon^{\top}\bar{x}_{K,i}\right)^{2}$;
       $\widehat{\textnormal{CI}}_{\upsilon}\leftarrow\left[\upsilon^{\top}\bar{x}_{K,i}-\widehat{\sigma}_{\upsilon}t_{1-\alpha/2,K-1}/\sqrt{K},\,\,\upsilon^{\top}\bar{x}_{K,i}+\widehat{\sigma}_{\upsilon}t_{1-\alpha/2,K-1}/\sqrt{K}\right]$
end for
Algorithm 1 Online Parallel Inference
Figure 1: Realizations of parallel computing and inference.
3 Theoretical guarantee
In this section, we provide a theoretical foundation for the confidence interval (10) constructed using the $t$-distribution. Recall that we consider a high level of confidence where the noncoverage level $\alpha$ can be potentially very small or decrease with the total sample size (or dimension). This level of validation requires a more stringent guarantee than just showing that
$\mathbb{P}(\upsilon^{\top}x^{*}\in\widehat{\textnormal{CI}})-(1-\alpha)\rightarrow 0,$ (11)
which can be derived from the convergence of relevant statistics in distribution as shown in other works. Our focus is to establish the bound of the relative error of coverage
$\Delta_{N}:=\sup_{\alpha(N)\leq\alpha<1}\left|\frac{\mathbb{P}(\upsilon^{\top}x^{*}\in\widehat{\textnormal{CI}})-(1-\alpha)}{\alpha}\right|,$
where $\alpha(N)$ goes to zero at an appropriate rate. Compared with (11), this bound offers a more rigorous assessment. It is critical in cases where we require high precision in our confidence assessments, ensuring that the constructed interval genuinely reflects the desired confidence level. For example, suppose we use Bonferroni method to construct simultaneous confidence intervals for $m$ parameters at overall level $0.95$ with large $m$, then the CI for each individual parameter should be at level $1-0.05/m$. In this case $\alpha=0.05/m$ and a small $\Delta_{N}$ is needed, while (11) is not sufficient. Also, it is important to make the dependence on level $\alpha$ explicit since we may consider a decreasing $\alpha$.
To derive the upper bound of the relative error, it is important to obtain the rate of convergence of the $t$-statistic. In the rest of this section, we will first explore the application of ASGD in each parallel run and then extend our results to a broader class of stochastic algorithms that meet certain mild assumptions.
3.1 Convergence characterization for ASGD
Among various stochastic approximation algorithms, SGD is notably convenient and popular. Its variant, ASGD, is also widely used and has been the subject of extensive study. Beyond the well-known asymptotic normality results related to convergence in distribution, the rate of convergence to normality is of growing interest and has been studied in the literature. Notably, [1] derived the non-asymptotic rate of convergence to normal using non-asymptotic rates of the martingale Central Limit Theorem (CLT), and [32] established a Berry–Esseen type bound for the Kolmogorov distance between the cumulative distribution functions of the ASGD estimator and its Gaussian analogue.
In this subsection, to better characterize the distributional approximation of the ASGD estimator, we develop a new Gaussian approximation of which the asymptotic normality is a direct consequence. Before presenting the main approximation result, we first introduce some regularity assumptions on the objective function and basic definitions. For any vector $\nu=(\nu_{1},\ldots,\nu_{m})^{\top}\in\mathbb{R}^{m}$, we use $|\nu|=\sqrt{\sum_{\ell=1}^{m}\nu_{\ell}^{2}}$ to denote its Euclidean norm.
Assumption 1.
There exist positive constants $\tau$ and $L$ such that
$(x-x^{\prime})^{\top}(\nabla F(x)-\nabla F(x^{\prime}))$ $\geq\tau|x-x^{\prime}|^{2},$
$|\nabla F(x)-\nabla F(x^{\prime})|$ $\leq L|x-x^{\prime}|.$
Assumption 2.
Denote $\Delta(x,\xi)=\nabla F(x)-\nabla f(x,\xi)$ for $x\in\mathbb{R}^{d}$ and $\xi\sim\Pi$. Given $q>4$, we have $\mathbb{E}_{\xi}|\Delta(x^{*},\xi)|^{q}<\infty$ and there exists some positive constant $\gamma$ such that for any $x,x^{\prime}\in\mathbb{R}^{d}$,
$\left(\mathbb{E}_{\xi}|\Delta(x,\xi)-\Delta(x^{\prime},\xi)|^{q}\right)^{1/q}\leq\gamma|x-x^{\prime}|.$
Assumption 3.
There exists some positive constant $\mathcal{L}$ such that for $x\in\mathbb{R}^{d}$,
$|\nabla F(x)-\nabla^{2}F(x^{*})(x-x^{*})|\leq\mathcal{L}|x-x^{*}|^{2}.$
Assumptions 13 are common and fairly mild in the context of convex optimization based on the SGD algorithm and its variants [4, 40]. For $n\geq 1$, we define
$\Gamma_{n}=\frac{1}{n}\sum_{k=1}^{n}U_{k}SU_{k}^{\top},\enspace\mathrm{where}\enspace U_{k}=\sum_{i=k}^{n}Y_{k}^{i}\eta_{k}$ (12)
with $Y_{k}^{k}={\bf{I}}_{d},Y_{k}^{i}=\prod_{l=k+1}^{i}({\bf{I}}_{d}-\eta_{l}\nabla^{2}F(x^{*})),i>k$. In the following theorem, we establish a Gaussian approximation result for the ASGD estimator.
Theorem 1.
Assume that $\{x_{i}\}_{i=1}^{n}$ is a SGD sequence defined by:
$x_{i}=x_{i-1}-\eta_{i}\nabla f(x_{i-1},\xi_{i}),\quad i=1,2,\ldots,$
where $\eta_{i}=\eta\times i^{-\beta}$ for some constant $\beta\in(1/2,1)$. Let $\bar{x}_{n}=n^{-1}\sum_{i=1}^{n}x_{i}$. Under Assumptions 13, on a sufficiently rich probability space, there exist a random vector $W_{n}\overset{\mathcal{D}}{=}\sqrt{n}(\bar{x}_{n}-x^{*})$ and a centered Gaussian random vector $Z_{n}\sim\mathcal{N}(0,\Gamma_{n})$, where $\Gamma_{n}$ is defined in (12), such that
$\mathbb{E}|W_{n}-Z_{n}|^{2}\lesssim\max\left(n^{1-2\beta},\frac{\log n}{n^{1-2/q}},\frac{|x_{0}-x^{*}|^{2}}{n}\right).$ (13)
Remark 3.
Theorem 1 reveals that the ASGD estimator can be approximated by a centered Gaussian random vector with approximately the same covariance matrix without imposing additional structural and moment assumptions and the approximation error is asymptotically negligible as long as $\beta>1/2$ and $|x_{0}-x^{*}|\ll n^{1/2}$. To the best of our knowledge, this is the first Gaussian approximation result for online estimators based on the SGD algorithm. It is worth noting the SGD iterates $x_{i}\in\mathbb{R}^{d},i=1,2,\ldots$, are neither independent nor stationary. Hence the existing strong invariance principle results for the partial sums of independent random elements (e.g., [15, 16, 5, 9]) or general stationary sequences (e.g., [38, 22, 2]) are not applicable here. To handle the nonstationary property of the sequence $\{x_{i}\}_{i\geq 1}$, we shall invoke the recently established strong approximation result for non-stationary time series ([25]). The detailed proof of Theorem 1 is given in the Appendix.
Remark 4.
It is worth noting the covariance matrix $\Gamma_{n}$ defined in (12) and hence the distribution of the coupled Gaussian vector $Z_{n}$ do not depend on the initial estimate $x_{0}$. Therefore, our result in  (13) can also be viewed as a quenched Gaussian approximation in the sense that the impact of the initial point $x_{0}$ diminishes, as asserted by the third term ${|x_{0}-x^{*}|^{2}}/{n}$ in the upper bound (13). As a direct consequence, the distribution of the random vector $\sqrt{n}(\bar{x}_{n}-x^{*})$ can be approximated by that of $\mathcal{N}(0,\Gamma_{n})$. Moreover, the multivariate central limit theorem (4) can be easily derived as $\Gamma_{n}$ converges to the sandwich form covariance matrix $\Sigma=A^{-1}SA^{-1}$; see, for example [27]. It is important to mention that our procedure does not rely on the convergence of $\Gamma_{n}$ to $\Sigma$, which can be slow and introduce additional approximation error in practical implementations. Our constructed $t$-statistic is asymptotically pivotal as long as (13) holds. Particularly, the simulation studies in Section 4 demonstrate that our procedure has better finite-sample performance than the oracle procedure based on the multivariate central limit theorem (4) with the population covariance matrix $\Sigma$ given.
The rate of convergence to normality plays a crucial role in assessing the approximation of the $t$-distribution. As we will discuss in a later section through a general theorem, the convergence of the $t$-statistic and the upper bound of the relative error relies on the convergence rate (of a single parallel run sequence) to normality.
3.2 Main results
As discussed above, there are many variants of SGD exhibiting asymptotic normality results, and the specific convergence rate to normality may vary across different algorithms. We will not study those rates for other algorithms rigorously since they are beyond the scope of this paper. To derive general results with the potential to be applied to different algorithms, we propose the following assumption.
Assumption 4 (Convergence rate to normality).
For a chosen stochastic algorithm and number of parallel runs $K$, let $\widehat{x}_{n}^{(k)}(k=1,...,K)$ denote the result at the $n$-th iteration of the $k$-th parallel run used in calculating the parallel average $\bar{x}_{K,n}$ in (8). There exists a centered Gaussian random vector $Z_{n}\sim\mathcal{N}(0,\Sigma_{n})$ (for some $\Sigma_{n}\in\mathbb{R}^{d\times d}$) such that
$\left(\mathbb{E}|\sqrt{n}(\widehat{x}_{n}^{(k)}-x^{*})-Z_{n})|^{2}\right)^{1/2}\lesssim\delta(n),$
where the approximation rate $\delta(n)\rightarrow 0$.
Theorem 1 demonstrates that if we employ ASGD as defined in (7), Assumption 4 is satisfied with $\Sigma_{n}=\Gamma_{n}$ defined in (12) and
$\delta(n)=\max\left(n^{1/2-\beta},\frac{\sqrt{\log n}}{n^{1/2-1/q}}\right)$
With this assumption, we are ready to show that the statistic in (9) is asymptotically pivotal with a specific convergence rate.
Theorem 2.
Suppose we run Algorithm 1 and Assumption 4 holds. For any $\upsilon$ and $\widehat{t}_{\upsilon}$ defined in (9) we have
$\sup_{z\in\mathbb{R}}\left|\mathbb{P}\left(\widehat{t}_{\upsilon}\geq z\right)-\mathbb{P}(T_{K-1}\geq z)\right|\lesssim(\delta(N/K))^{1/4},$
where $T_{K-1}$ is a random variable following $t$ distribution with degree of freedom $K-1$, $N$ is the total sample size and $K$ is the number of parallel runs. Consequently, for any confidence level $\alpha\in(0,1)$,
$\left|\frac{\mathbb{P}\left(|\widehat{t}_{\nu}|\geq t_{1-\alpha/2,K-1}\right)}{\alpha}-1\right|\lesssim\alpha^{-1}\delta(N/K)^{1/4},$
where $t_{1-\alpha/2,K-1}$ is the $(1-\alpha/2)\times 100\%$ percentile for the $t_{K-1}$ distribution and the constant in $\lesssim$ does not depend on $\alpha$. For $\alpha(N)$ goes to zero with $\delta(N/K)^{1/4}\ll\alpha(N)$, the relative error of coverage goes to zero when $\alpha\geq\alpha(N)$, i.e.,
$\Delta_{N}=\sup_{\alpha(N)\leq\alpha<1}\left|\frac{\mathbb{P}(\upsilon^{\top}x^{*}\in\widehat{\textnormal{CI}})-(1-\alpha)}{\alpha}\right|\rightarrow 0.$
The results suggest that any stochastic algorithm demonstrating appropriate convergence in certain scenarios can be selected, provided its single sequence exhibits convergence towards normality. The convergence of the $t$-statistic, as well as the relative error in the coverage of the confidence interval, can be bounded based on the rate of convergence to normality. Furthermore, this study comprehensively examines the reliance on the value of $\alpha$. The uniform convergence of $\Delta_{N}$ indicates that an extremely small $\alpha$, or decreasing $\alpha$, is feasible.
4 Experiment
4.1 Simulation
In this section, we investigate the empirical performance of our proposed parallel inference under the linear regression model and logistic regression model. The true coefficient of interest $x^{*}$ is a $d$-dimensional vector with $x^{*}=(0,1/d,2/d,...,(d-1)/d)$. We generate a sequence of i.i.d. random samples $\{(a_{i},b_{i})\}_{i=1}^{n}$, where $a_{i}$ stands for the explanatory variable generated from $\mathcal{N}(0,\bf{I}_{d})$, and $b_{i}$ stands for the response variable. In the linear regression model, we have
$b_{i}=a_{i}^{T}x^{*}+\epsilon_{i},$
where $\epsilon_{i}$ follows $\mathcal{N}(0,1)$ independently. The corresponding loss function is $f(x,\xi_{i})=(a_{i}^{T}x-b_{i})^{2}/2$. In the logistic regression model, $b_{i}\in\{0,1\}$ is generated from a Bernoulli distribution, where
$\mathbb{P}(b_{i}=1|a_{i})=\frac{1}{1+\exp(-a^{T}_{i}x^{*})},$
and the loss function is logit loss as $f(x,\xi_{i})=(1-b_{i})a_{i}^{T}x+\log(1+\exp(-a^{T}_{i}x))$. We employ ASGD for our parallel method, with $\beta=0.505$, and $\eta=0.5$, consistent with the settings used in [17, 40]. We consider the case of marginal inference of coordinates, that is the vector $\upsilon$ in linear functional is chosen as the canonical basis. To analyze the empirical performance, we record the coverage of the constructed confidence intervals, the relative error of coverage $\Delta_{\alpha}$ as defined in (3), the length of the confidence intervals, and the running time. All reported results are the average of 10000 independent trials.
Figure 2: Effect of $K$. Plot (a): relative error of coverage; plot (b): the length of confidence interval. The nominal coverage probability is 0.99. The total sample size $N$ is $60000$ for linear models and $200000$ for logistic models.
Choice of K. We first examine the effect of $K$. We construct $99\%$ confidence intervals ($\alpha=0.01$) with a total sample size of $60000$ ($N=nK=60000$) for linear regression, and with a total sample size of $200000$ ($N=nK=200000$) for logistic regression. From Figure 2, we observe that a small $K$ may decrease the bias of coverage in some challenging scenarios, such as logistic regression with $d=20$. On the other side, it will result in longer confidence intervals. However, when $K$ falls within a reasonable range, say between $2$ to $11$, the results appear satisfactory and are not overly sensitive to the choice of $K$. In the following simulation results, we will use $K=6$.
Compare to another method. We compare the finite sample performance of our proposed inference method, referred to as the parallel method, with that of the state-of-the-art method: the random scaling method [17], which also leverages an asymptotic pivotal statistic. The confidence interval constructed by the random scaling method is given in (5), and we obtain critical values through Monte Carlo simulation as tabulated in Table 3. We did not include comparisons with other methods such as the Plug-in [4] or Online Batch-means [40], as the random scaling method has already demonstrated comparable coverage to the Plug-in method, superior coverage compared to Online Batch-means, and faster computing times. For both methods, we apply the ASGD algorithm with $\beta=0.505$ and $\eta=0.5$. The number of parallel runs, $K$, is set to $6$ for the parallel method. We consider constructing confidence intervals every 600 samples. Overall, the performance of the parallel method is satisfactory and better than that of the random scaling method, with faster convergence, comparable confidence interval lengths, and less computation.
(a) $\alpha=0.05$
(b) $\alpha=0.01$
(c) $\alpha=0.001$
Figure 3: Linear Regression $d=20$: Left: relative error of coverage; Middle: empirical coverage; Right: length of confidence intervals.
(a) $\alpha=0.05$
(b) $\alpha=0.01$
(c) $\alpha=0.001$
Figure 4: Logistic Regression $d=20$: Left: relative error of coverage; Middle: empirical coverage; Right: length of confidence intervals.
(a) Linear regression
(b) Logistic regression
Figure 5: Computation time: d = 20
(99.9% CI) Coverage (Relative error) Length (std) Time
Linear d=5 Parallel 0.9991 (0.080) 0.055 (0.00797) 0.918
Random Scale 0.9986 (0.420) 0.046 (0.00840) 2.262
Linear d=20 Parallel 0.9990 (0.015) 0.058 (0.00563) 1.135
Random Scale 0.9987 (0.310) 0.046 (0.00431) 3.025
Logistic d=5 Parallel 0.9990 (0.000) 0.070 (0.01023) 2.861
Random Scale 0.9976 (1.380) 0.057 (0.01120) 7.044
Logistic d=20 Parallel 0.9991 (0.064) 0.088 (0.00689) 3.363
Random Scale 0.9960 (2.984) 0.069 (0.00738) 9.384
(99% CI) Coverage (Relative error) Length (std) Time
Linear d=5 Parallel 0.9896 (0.042) 0.032 (0.00467) 0.918
Random Scale 0.9885 (0.146) 0.031 (0.00572) 2.262
Linear d=20 Parallel 0.9903 (0.036) 0.034 (0.00340) 1.135
Random Scale 0.9888 (0.124) 0.032 (0.00293) 3.025
Logistic d=5 Parallel 0.9889 (0.106) 0.041 (0.00601) 2.861
Random Scale 0.9840 (0.598) 0.039 (0.00762) 7.044
Logistic d=20 Parallel 0.9898 (0.017) 0.051 (0.00405) 3.363
Random Scale 0.9784 (1.156) 0.047 (0.00502) 9.384
(95% CI) Coverage (Relative error) Length (std) Time
Linear d=5 Parallel 0.9503 (0.006) 0.021 (0.00300) 0.918
Random Scale 0.9454 (0.090) 0.021 (0.00384) 2.262
Linear d=20 Parallel 0.9509 (0.017) 0.022 (0.00218) 1.135
Random Scale 0.9461 (0.078) 0.021 (0.00197) 3.025
Logistic d=5 Parallel 0.9477 (0.046) 0.026 (0.00383) 2.861
Random Scale 0.9363 (0.274) 0.026 (0.00512) 7.044
Logistic d=20 Parallel 0.9499 (0.024) 0.033 (0.00258) 3.363
Random Scale 0.9222 (0.554) 0.032 (0.00337) 9.384
Table 2: Compare parallel method and random scaling method: $\alpha=\{0.001,0.01,0.05\}$ ($99.9\%$, $99\%$ and $95\%$ confidence). Total sample size $N=60000$ for Linear regression, and $N=200000$ for Logistic regression. (inference is performed every 600 samples)
In Figures 3 and 4, we present results for confidence intervals where the nominal coverage probability is set at $0.95$, $0.99$, and $0.999$, i.e., $\alpha={0.05,0.01,0.001}$ for both linear regression and logistic regression. We plot the relative error of coverage, the empirical coverage rate, and the length of the confidence intervals. We also compare the running time of a single trial in Figure 5. More results are summarized in Table 2 and Appendix. The relative error of our parallel method converges to zero faster than that of the random scaling method in all cases. The advantage becomes more obvious as confidence levels increase. Note that in logistic regression with $d=20$, both methods exhibit relatively large errors when the sample size is small. In this case, the parallel method converges more slowly, which can be attributed to the fact that data splitting in the parallel run exacerbates the issue of a small sample size. However, as the sample size increases, the convergence rate of the parallel method improves and eventually surpasses that of the random scaling method. The lengths of the confidence intervals are comparable between the two methods, with those derived from the parallel method being slightly larger. We also observe that our parallel method has a distinct advantage in terms of computing time, as it does not necessitate additional computations at each iteration, such as updating a $d$ by $d$ matrix, which is required by the random scaling method. Apart from the SGD update, the only additional computation needed for inference is calculating a sample covariance matrix or a sample variance (for the linear functional). This computation is minimal, making the inference process almost cost-free. The advantage in computing becomes even more significant when utilizing parallel computing across different cores.
Compare to oracle. We also compare our method to the oracle approach, which constructs confidence intervals using the true limiting covariance matrix $\Sigma$ and the principles of asymptotic normality. The oracle method is given by:
$\widehat{\textnormal{CI}}_{n,\textnormal{oracle}}=\left[\upsilon^{\top}\bar{x}_{n}-z_{1-\alpha/2}\sqrt{\frac{\upsilon^{\top}\Sigma\upsilon}{n}},\ \upsilon^{\top}\bar{x}_{n}+z_{1-\alpha/2}\sqrt{\frac{\upsilon^{\top}\Sigma\upsilon}{n}}\right].$
In the linear regression model described above, the limiting covariance matrix $\Sigma={\bf{I}}_{d}$. We focus on linear regression here since the true limiting covariance matrix is straightforward to compute. As illustrated in Figure 3 the coverage achieved by the parallel method surpasses that of the oracle method. This could be attributed to a discrepancy between the finite sample covariance matrix of ASGD and the limiting covariance $\Sigma$. Employing an asymptotic pivotal statistic helps to mitigate the impact of this difference. For further details and discussion on this topic, refer to Section 3.1.
4.2 Hand-written digit dataset
To further explore the application of confidence intervals, we consider the task of estimating the mean image for each digit in the MNIST handwritten digit dataset.
The MNIST dataset comprises $60000$ training images, each each measuring $28\times 28$ pixels in dimension and labeled with digits ranging from 00 and $9$. For each label $m=0,1,\dots,9$, we hypothesize the existence of a mean image $x^{*,m}\in\mathbb{R}^{d},d=28\times 28$, and individual image instances $\{\xi_{i}^{(m)}\}$ sampled from a normal distribution with mean $x^{*,m}$ and an unknown variance. Our goal is to estimate these mean images. The objective function is defined as $f(x,\xi)=\frac{1}{2}\|\xi-x\|^{2}$. We employ the online parallel algorithm (1) with $K=6$ and ASGD with a step size at the $n$-th iteration $\eta_{n}=n^{-0.505}$. The parallel means computed from this process serve as our final mean estimates. We construct coordinate-wise confidence intervals based on Algorithm 1, choosing vectors $\upsilon=e_{j},j=1,...,784$. We visualize the mean images, noting that in the grayscale representation, 00 denotes gray, $-1$ denotes black, and $1$ denotes white. Our approach further includes denoising; that is, truncating (or ‘shrinking’) values below a certain threshold to $-1$ to make the mean image sharper. Traditionally, this threshold lacks formal guidance. We first try a one-size-fits-all threshold. In Figure 6 (a) , the uniform threshold is set at 00. The results are not satisfactory; significant portions of the digit $5$ were missing, and it did not make sense to shrink the upper part of $6$. Changing the uniform threshold to a smaller number, $-0.5$, as shown in Figure 6 (b), the denoising step has no effect. It remains unclear which threshold is effective, and it is uncertain if a single, uniform threshold is appropriate. Then, we refined the denoising step by leveraging confidence intervals constructed by the parallel method: a mean value is set to 00 if its coordinate-wise upper confidence bound is below 00. Given the desire to preserve sufficient pixel detail, we opt for a high confidence level, adjusted for the number of parameters. Thus, we set our confidence level to $1-\alpha/d$ with $\alpha=0.001$, effectively achieving a $99.9\%$ confidence interval across all coordinates simultaneously. The results in Figure 6 (d) demonstrate that using confidence intervals yields more accurate and visually coherent mean images.
(a) Original mean
(b) $t=0$
(c) $t=-0.5$
(d) Adaptive
Figure 6: Mean image before and after denoising. (a) shows the original estimated mean before denoising; (b) uses a uniform threshold $t=0$; (c) uses a uniform threshold $t=-0.5$; (d) applies an adaptive threshold based on the upper bound of the confidence interval.
5 Discussion
In this paper, we introduce a novel inference framework designed to construct confidence intervals for model parameters by employing stochastic algorithms in an online environment. This method stands out for its simplicity and ease of implementation, offering flexibility across various algorithms. The construction of confidence intervals is computationally efficient and almost free (post-SGD update). Furthermore, we bolster our approach with rigorous theoretical guarantees, demonstrating its capability to facilitate inference at a high confidence level.
Appendix A Additional numerical results
In Table 3, we provide critical values used in the random scaling method. In Figure 7, 8, 9 and Table 2, we provide results for linear regression and logistic regression when $d=5$ with same settings as described in Section 4.
Probability 97.5% 99.5% 99.95%
Critical Value 6.474 10.0544 14.76972
Table 3: Asymptotic one-sided critical values for asymptotic pivotal statistic in the random scaling method (5) via Monte Carlo simulation with 1,000,000 samples.
(a) $\alpha=0.05$
(b) $\alpha=0.01$
(c) $\alpha=0.001$
Figure 7: Linear Regression $d=5$: Left: relative error of coverage; Middle: empirical coverage; Right: length of confidence intervals.
(a) $\alpha=0.05$
(b) $\alpha=0.01$
(c) $\alpha=0.001$
Figure 8: Logistic Regression $d=5$: Left: relative error of coverage; Middle: empirical coverage; Right: length of confidence intervals.
(a) Linear, $d=5$
(b) Logistic, $d=5$
Figure 9: Computation time: d = 5